The heat of combustion of liquid cumene (C9H12) is -5215.5 kJ/mol at 1 atm and 298 K

a) What is the energy of combustion at 1 atm and 298 K?
deltacap.gifEcomb =  kJ/mol

b) Use Thermodynamic Properties to determine its heat of formation.
deltacap.gifHf =  kJ/mol
Answer #1

cas Cg Hiz, et 12 02 g CO2 + 6H₂O x ang= g -12=-3 DH = AE + Ang RT. - AE- AH - A và RT = -5215.5 kJ 43x 3X 8.3) 4 5 15.5 kJ